# Data-driven computation of adjoint sensitivities without adjoint solvers: An application to thermoacoustics

### v1

@@ IX Bifurcations to nonlinear regimes
Figure 12: Computation of chaotic adjoint sensitivities using thermoacoustic ESN. The adjoint sensitivity for a single trajectory at $(\beta=7.6,\;\tau=0.22)$ is shown on the left. The probability distribution of the sensitivity is computed from 2000 short trajectories with integration time of 0.5 LT (c,d) and 1.0 LT (e,f).
X Conclusion
@@
Adjoint methods are established tools for gradient-based optimization. Although adjoint methods offer an accurate and efficient gradient computation, they are code specific. This means that an adjoint solver needs to be derived/implemented for new problems of interest. In this work, we propose a strategy to infer adjoint sensitivities from data (observables). We showcase the method on a prototypical nonlinear time-delayed thermoacoustic system. This system is described by a nonlinear wave equation. The parameters of interest are the heat release strength and the time-delay at the heat source location. By varying these parameters, the system displays fixed-point, limit-cycle, quasiperiodic, and chaotic solutions. The cost functional is the time-average acoustic energy, which we wish to minimize to suppress an instability.
@@
First, we propose a parameter-aware Echo State Network (ESN). This network learns the parameterized dynamics, i.e., it infers how the dynamics change with varying parameters, thereby making predictions for parameters that are not used during training. Second, we derive the adjoint parameter-aware Echo State Network, which is used to compute the adjoint sensitivities of the physical system. Third, we design a parameter-aware ESN motivated by the physics of thermoacoustics (thermoacoustic ESN, or T-ESN in short) by (i) making the ESN explicitly time-delayed, and (ii) designing the input matrix in consideration of acoustic physics. We extend the derivation of the adjoint to the time-delayed case. We find that the T-ESN accurately predicts the dynamics of different nonlinear regimes, and successfully infers the adjoint sensitivities. These results are validated against the ground truth generated from integration of the direct and adjoint equations of the original system. We optimize the acoustic energy with steepest gradient-descent, in which the gradients are computed by only one T-ESN. Once the T-ESN has been trained, the algorithm runs autonomously, i.e., we do not run any additional simulations of the original model. As the optimization proceeds with the iterations, the acoustic energy and the amplitude of the oscillations decrease, finally reaching a fixed-point solution with zero acoustic energy. Fourth, we investigate the robustness to noisy data and chaotic regimes. The T-ESN can qualitatively predict bifurcations to nonlinear regimes beyond the training range of parameter values. Quantitative accuracy increases when the range of training parameters covers the regime in which we are interested. This work opens up possibilities for data-driven optimization without developing code-specific adjoint solvers.

### v2

@@ IX Bifurcations to nonlinear regimes
Figure 15: Computation of chaotic adjoint sensitivities using thermoacoustic Echo State Network. The adjoint sensitivities for a single trajectory at $(\beta=7.6,\;\tau=0.22)$ is shown on the left (a,b). The probability distribution of the sensitivity is computed from 2000 short trajectories with integration time of 0.5 Lyapunov Time (c,d) and 1.0 Lyapunov Time (e,f).
X Computation of sensitivities to initial conditions
The adjoint method allows us to efficiently compute sensitivities to initial conditions. In this section, we infer the sensitivity to initial conditions, i.e., $\mathbf{y}(0)=[\mathbf{\bm{\eta}}(0);\;\mathbf{\bm{\mu}}(0)]$, using the T-ESNs trained in the previous sections. The objective functional is the acoustic energy of the final flow state. The sensitivities to perturbations to the first velocity and pressure modes, $\eta_{1}$ and $\mu_{1}$, are shown in Fig. 16 for a trajectory from exemplary limit cycle $(\beta=4.5,\;\tau=0.12)$, quasiperiodic $(\beta=6.1,\;\tau=0.2)$, and chaotic $(\beta=7.6,\;\tau=0.22)$ regimes. To assess the accuracy of the estimates, we repeat this procedure for multiple independent trajectories. Because in gradient-based optimization we are interested in the gradient direction, we compute the angle between the sensitivity gradients of the original system and the T-ESN estimates as an indicator of the accuracy. The mean of the computed angles over 20 trajectories and 5 realisations of T-ESN are provided in Tab. 3 for different integration times. We find that T-ESN achieves high accuracy for all integration times in the limit cycle regimes and for short integration times in quasiperiodic and chaotic regimes.
Figure 16: Computation of adjoint sensitivities to initial conditions using thermoacoustic Echo State Network. The adjoint sensitivity for a single trajectory at unseen (a,b) limit cycle $(\beta=4.5,\;\tau=0.12)$, (c,d) quasiperiodic $(\beta=6.1,\;\tau=0.2)$, (e,f) chaotic $(\beta=7.6,\;\tau=0.22)$ regimes. The sensitivity of the acoustic energy of the final flow state to perturbation to the first velocity mode $\eta_{1}$ (top row) and to the first pressure mode $\mu_{1}$ (bottow row) are shown for increasing integration times.
Table 3: Angle between the sensitivity gradients of the original system to initial conditions and the estimates from thermoacoustic Echo State Network (T-ESN). The mean angle over 20 independent trajectories and 5 realisations of T-ESN is given in degrees.
Limit cycle Quasiperiodic
$\;(\beta=4.5,\;\tau=0.12)\;$ $\;(\beta=6.1,\;\tau=0.2)\;$
Integration Time 2.0 $2.2$ $6.8$
4.0 $2.3$ $12.8$
10.0 $3.2$ $16.2$
Chaotic
$\;(\beta=7.6,\;\tau=0.22)\;$
0.5 LT $8.0$
1.0 LT $15.5$
2.0 LT $34.3$
XI Discussion
In this section, we offer a comparison of the proposed ESN-based adjoint solver with traditional, i.e., physics-based, adjoint solvers for gradient-based optimization. With the physics-based approach, we need (i) governing equations (plus boundary conditions and initial conditions), (ii) a numerical scheme to solve them, (iii) the linearization (tangent model), and (iv) its dual with respect to a chosen inner product (adjoint code). We need to derive/modify the adjoint code any time that any of the above is modified. On the other hand, in the ESN-based approach the adjoint code is derived once and does not change, i.e., the adjoint equations of the ESN described by Eqs. (13) and (19) are independent of the data and system under investigation. Similarly, the adjoint equations of the time-delayed ESN described by Eq. 28 are applicable to any system with a constant time-delay. The ESN-based approach is also particularly attractive when we have experimental data, in which the equations might be unknown (or partly unknown) and the data contains noise, because it learns the dynamics from observables.
As a data-driven approach, the ESN-based method relies on (i) having sufficient data that covers the solution space of the problem, and (ii) training and validation of the ESN. When utilizing a parameter-aware ESN for the computation of parameter sensitivities, as the number of system’s parameters increase, more training regimes are required. This is in contrast with physics-based adjoint solvers, which are derived from the system’s equations that generalise over the parameter space. Thus, if the ESN-based adjoint solver is sought to replace a physics-based adjoint solver, the advantage can vary from case to case depending on the cost of data acquisition and simulation time. Because the network learns a continuous parameterization of the dynamics even if it is trained on data from a discrete set of parameters, the adjoint can be computed around any set of parameters. (With a continuous parameterization of the objective functional, a future research direction is to combine Bayesian optimisation [e.g., 7], which relies on sampling via the ESN, and adjoint-based ESN gradients.) The physics-based adjoint solver depends on the modelling assumptions, whereas the accuracy of the ESN-based adjoint solver depends on data and the hyperparameters. In high-dimensional systems, the ESN-based adjoint solver can be combined with a data-driven model reduction techniques, e.g., autoencoding as in [26].
Finally, regarding computational costs, the training of the ESN is faster compared to other neural networks that are trained with stochastic gradient descent and backpropagation, and the T-ESN runs faster than the original solver when simulating the thermoacoustic system considered in this paper. Although this can be case dependent, in general the propagation of the ESN equations is fast thanks to the sparse weight matrices even though the reservoir state can be high dimensional (here, 1200). An optimization of the code for matrix multiplications on a GPU, e.g., using JAX [49], is scope for future work.
XII Conclusion
@@
Adjoint methods are established tools for gradient-based optimization. Although adjoint methods offer an accurate and efficient gradient computation, they are problem specific. In this work, we propose a strategy to infer adjoint sensitivities from data (observables). We showcase the method on a prototypical nonlinear time-delayed thermoacoustic system. This system is described by a nonlinear wave equation. The parameters of interest are the heat release strength and the time-delay at the heat source location. By varying these parameters, the system displays fixed-point, limit-cycle, quasiperiodic, and chaotic solutions. The cost functional is the time-averaged acoustic energy, which we wish to minimize to suppress an instability.
@@
First, we propose a parameter-aware Echo State Network (ESN). This network learns the parameterized dynamics, i.e., it infers how the dynamics change with varying parameters, thereby making predictions for parameters that are not used during training. Second, we derive the adjoint parameter-aware Echo State Network, which is used to compute the adjoint sensitivities of the physical system. Although the proposed framework is general and has been applied to prototypical low-dimensional dynamical systems without any modification of the standard architecture, we find that for the thermoacoustic system of interest embedding physical knowledge improves generalizability and requires less data. Therefore, third, we design a parameter-aware ESN motivated by the physics of thermoacoustics (thermoacoustic ESN, or T-ESN in short) by (i) making the ESN explicitly time-delayed, and (ii) designing the input matrix in consideration of acoustic physics. We extend the derivation of the adjoint to the time-delayed case. We find that the T-ESN accurately predicts the dynamics of different nonlinear regimes, and successfully infers the adjoint sensitivities. These results are validated against the ground truth generated from integration of the direct and adjoint equations of the original system. We optimize the acoustic energy with steepest gradient-descent, in which the gradients are computed by only one T-ESN. Once the T-ESN has been trained, the algorithm runs autonomously, i.e., we do not run any additional simulations of the original model. As the optimization proceeds with the iterations, the acoustic energy and the amplitude of the oscillations decrease, finally reaching a fixed-point solution with zero acoustic energy. Fourth, we investigate the robustness to noisy data and chaotic regimes. The T-ESN can qualitatively predict bifurcations to nonlinear regimes beyond the training range of parameter values. Quantitative accuracy increases when the range of training parameters covers the regime in which we are interested. Finally, in addition to parameter sensitivities, we infer sensitivities to initial conditions. This work opens up possibilities for data-driven gradient-based optimization without developing code-specific adjoint solvers.